# CL-RAG: Bridging the Gap in Retrieval-Augmented Generation with Curriculum Learning

### v1

@@ 3.3 Curriculum Learning for Retriever
$\mathcal{D}_{rerank}=\{d_{1},d_{2},...,d_{n}\}.$
@@
Next, we construct data with different difficulty levels for curriculum learning, aiming to first train the retriever to distinguish documents with obvious quality gaps and then train it to distinguish documents with slightly quality differences. Specifically, we divide $\mathcal{D}_{rerank}$ into three groups $\{\mathcal{D}_{1},\mathcal{D}_{2},\mathcal{D}_{3}\}$ based on the ranking order from RALM. These groups contain documents in the ranges $[1,n_{1}]$, $[n_{1},n_{2}]$, and $[n_{2},n]$, respectively, representing good documents, sub-optimal documents, and hard negative documents. During training stages, we sample $k_{1}$, $k_{2}$, and $k_{3}$ documents from each of the three groups, while $k_{1}+k_{2}+k_{3}=k$, and gradually increase $k_{1}$ and decrease $k_{2}$ and $k_{3}$, narrowing the quality gap between the sampled documents. We have also designed a tiered loss function to better fit our approach:
$\mathcal{L}(q,\mathcal{D}_{k})=\\ -\sum_{d_{i,j}\in{\mathcal{D}_{k}}}\sum_{j>i}\frac{j-i}{n-1}log(\frac{e^{s_{i}}}{e^{s_{i}}+e^{s_{j}}}),$
@@
where $\mathcal{D}_{k}$ denotes a subset of $k$ documents selected from $\mathcal{D}$, $n$ represents the total number of documents involved in the ranking, and $i$,$j$ are the ranks of the documents, $s_{i}$ denotes the similarity score between the query $q$ and the document $d_{i}$. We start by distinguishing documents with obvious quality differences and gradually move to distinguishing documents with slight quality differences. This stage-by-stage approach helps bridge the gap between the original retriever and the retriever aligned with RALM preferences.
4 Experiments
4.1 Experimental Setting
Method NQ TriviaQA PopQA HotpotQA Avg.
EM F1 EM F1 EM F1 EM F1 EM F1
$Llama3_{8B}$ 39.51 53.37 72.74 81.80 46.07 52.16 25.23 39.46 45.89 56.70
$With$ $RALM$ $Training$
$RAAT$ 31.65 40.39 71.84 77.02 38.49 41.87 22.16 30.26 41.04 47.39
$RALM_{golden}$ 42.90 52.30 79.55 83.84 50.06 52.45 31.25 41.96 50.94 57.64
$RALM_{top5}$ 48.15 58.95 80.32 85.05 57.58 60.13 36.57 48.46 55.66 63.15
$RetRobust$ 47.83 58.60 81.25 85.96 57.51 60.31 37.02 48.99 55.90 63.47
$With$ $Retriever$ $Training$
$RA$-$DIT$ 49.52 59.94 76.62 81.10 61.73 64.47 36.14 47.82 56.00 63.33
$Replug$ 52.52 62.28 76.86 81.59 63.75 66.35 37.75 49.22 57.72 64.86
$LLM$-$Emebedder$ 51.85 61.41 82.60 87.20 64.24 66.69 38.80 50.51 59.37 66.45
$CL$-$RAG(Ours)$ 53.01 62.51 82.97 87.57 66.63 69.12 39.11 51.19 60.43 67.60
Table 1: Experimental results for EM and F1 scores(%) on four open-domain QA datasets compared with multiple baselines in RALM or Retriever training. "$RALM_{CL}$" refers to the RALM trained with CL strategy.
4.1.1 Datasets
Open-domain QA
We considered four open-domain QA datasets: single-hop question datasets Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and PopQA (Mallen et al., 2023), as well as the multi-hop question dataset HotpotQA (Yang et al., 2018). For each dataset, we employed Contriever-MSMARCO (Izacard et al., 2021) as the retriever. Following previous settings (Asai et al., 2024), for NQ, TriviaQA, and HotpotQA, we retrieved the corresponding documents from the 2018 Wikipedia corpus provided by Izacard et al. (2023). For PopQA, we retrieved the documents from the 2020 Wikipedia corpus. Noting that PopQA only contains a test set, which we split into two non-overlapping parts for training and testing, respectively. We utilized datasets from the KILT benchmark (Petroni et al., 2021), more detailed information about the datasets is provided in Appendix A.
Robustness Test
4.1.2 Evaluation Metrics
We employ Exact Match (EM) and F1 score as evaluation metrics. Specifically, EM assesses whether the model-generated answers are identical to the correct answers. Meanwhile, the F1 score integrates Precision and Recall to measure the accuracy and coverage of the model in generating answers.
4.1.3 Baselines
We utilized several training methods for RALM and Retriever as baselines, and firstly evaluated the performance of all the RALM training methods: (1)RAAT: Fang et al. (2024) enhances the robustness of RALMs through adversarial training. (2)RALMtop5: The most basic approach in RALM training by directly employing the $top$-$5$ retrieved documents. (3)RALMgolden: Lin et al. (2023) added the golden document to the retrieved documents to train the RALM. (4)RetRobust: Yoran et al. (2023) randomly injected irrelevant and counterfactual documents into the retrieved documents to expose the model to diverse types of noise during training.
Then, we continued to assess the performance of retrievers on the best-performing RALM: (1)Replug: Shi et al. (2023) minimized the KL divergence between the score of retriever’s and the model’s preference distribution to train a retriever that better aligns with the LLM. (2)RA-DIT: Lin et al. (2023) combined the method of finetuning the retriever using KL divergence and finetuning the RALM with retrieved documents that include golden documents as context. (3)LLM-Embedder: Zhang et al. (2024) defines the quality of document $d$ based on how much it improves the correct answer’s ranking in LLM’s response and proposes a fine-grained hierarchical loss function for training.
4.1.4 Implementation Details
RLAM Training
We trained our model based on LLaMA3-8B-Instruct (Dubey et al., 2024), using five retrieved documents. We randomly selected 1500 samples from the training sets of four datasets, totaling 6000 samples, and conducted LoRA finetuning based on LLaMA Factory (Zheng et al., 2024), with the LoRA rank set to 8, a learning rate of 1e-4, a gradient accumulation step of 8, and a warmup ratio of 0.1. All experiments were conducted on a A800 80G GPU card.
Retriever Training
We trained our model based on Contriever-MSMARCO (Izacard et al., 2021). By randomly sampling from the training sets of four datasets, we constructed a retriever training set comprising 100,000 samples. We set the number of input retrieved documents $k$ to 5, the learning rate of 1e-5, the number of epochs to 3, and the batch size to 64. For our three-stage curriculum learning settings, the number of reranked documents $n$ is set to 20. $n_{1}=[1,3,5]$ for each training epoch and $n_{2}$ is set consistently to 15. The sampling parameters for each stage are $k_{1}=[1,3,5]$, $k_{2}=[2,2,0]$, and $k_{3}=[2,0,0]$.
4.2 Main Results
Method Irrelevant Counterfactual
EM F1 EM F1
$Llama3_{8B}$ 44.10 54.79 44.32 54.55
$RAAT$ 39.17 45.38 39.27 45.65
$RALM_{golden}$ 49.36 56.07 48.25 54.73
$RALM_{top5}$ 53.48 61.05 53.63 61.13
$RetRobust$ 53.85 61.43 54.16 61.77
$RALM_{CL}$ 56.17 63.42 56.21 63.47
Table 2: Experimental results on robustness test(%) over irrelevant and counterfactual documents. We report the average results on each dataset here, and the complete results are provided in Appendix C.
The overall results are shown in Table 1, where the underlined and boldface items indicate the best results under each setting. We notify that the training of all methods adhered to a unified model and settings, and were tested under a 0-shot setting, with the prompts provided in the Appendix B.
With RALM Training
The results in the upper half of Table 1 demonstrate that our training method achieves better performance than previous methods across all test sets, thereby validating the effectiveness of our approach. It is noted that the model trained with the RAAT method performs worse than the baseline model. This is because its training data follows the setting of $\{d_{golden}\oplus d_{noise}\}$. While this setting trains the model to be robust against noise, the inclusion of golden documents in all data deviates from the conventional RAG setup, resulting in poorer performance under the standard setting. Training using documents that contain a golden document can enhance the model’s basic performance, i.e., its ability to extract correct answers from documents. However, due to the limitations of the retriever, the retrieved documents may not always contain the necessary information to answer the question. Therefore, training solely for answer extraction is insufficient. It is also necessary to further train the model to infer answers from more complex documents. RetRobust considers a more complex setting by accounting for both irrelevant document noise and counterfactual document noise. However, it neglects the incorporation of golden documents, which are beneficial in the early stages of model training. In contrast, our proposed method considers all types of document combinations, thereby enabling the model to adapt to the diverse documents generated by retrieval. Therefore, it achieves the best performance, with improvements of 2.18% in the average EM score and 1.85% in the average F1 score.
With Retriever Training
In the settings that incorporate retriever training, our method also achieves the best average performance. As shown in the lower half of Table 1, RA-DIT utilizes the model from its experimental setting, namely $RALM_{top5}$, while the other methods all employ the best-performing $RALM_{CL}$. We found that on datasets where LLM itself performs well, such as TriviaQA, the gains from further improving the retriever’s performance are minimal. However, for datasets like PopQA, which involve questions about specific entities, enhancing the retriever enables it to more accurately fetch entity-related documents, resulting in a significant boost in performance.
4.3 Robustness Test
To demonstrate the superiority of our RALM in combating noise, we also conducted tests under document settings that included noise. The results are shown in Table 2. Under the influence of document noise, the performance of all models declined to some extent. The impact of irrelevant document noise was comparable to that of counterfactual noise. The experimental results show that our model maintained its leading performance even when different types of noise were added.
4.4 Further Study of Retriever Training
Method Avg.
R@1 R@5 R@10
$Baseline$ 44.35 67.40 74.26
$Replug$ 49.10 72.32 78.43
$w/o\ RALM$ 43.82 67.25 74.25
$LLM$-$Embedder$ 49.45 71.48 77.78
$w/o\ RALM$ 47.78 69.78 76.29
$CL$-$RAG$ 50.40 73.29 79.55
$w/o\ RALM$ 48.55 70.71 77.33
Table 3: Experimental results on Recall rates(%) for retrievers trained with different methods. "Baseline" refers to the retrieval results of the raw retriever and "w/o RALM" indicates using the raw LLM without employing the well-trained RALM. Detailed results for each dataset are provided in Appendix D.
To demonstrate that using the trained RALM to guide the finetuning of the retriever is superior, we compared the retrieval results of the retriever finetuned with the raw LLM preferences to those finetuned with the preferences of RALM. We reported the recall rate of correct answers in the retrieved documents in Table 3. The experimental results show that finetuning with the preferences of RALM generally outperforms finetuning with the raw LLM preferences. We observed that Replug, which assesses document quality based on answer probability, led to a significant drop in recall rate. This is because when the LLM itself is capable of providing the correct answer, using answer probability as a criterion to judge the quality of documents becomes unstable. In contrast, the trained RALM has a higher ability to discern documents and can more accurately differentiate between good and bad ones. The experiments on final answer generation are provided in Appendix D. It is worth emphasizing that even when the same trained RALM is used for guidance, our method still maintains a leading position.
4.5 Ablation Study
stage3stage2stage154545454565656565858585860606060w/o CLw/o CLEM(%)With RALM TrainingWith Retriever Training
Figure 4: Ablation studies for our CL-RAG framework(%). We report the average EM scores and the complete results are provided in Appendix E.
To systematically elucidate the contribution of each training stage to the overall performance, we conducted an ablation study. Specifically, we evaluated the intermediate models generated at each stage of our CL-RAG framework in a stage-by-stage manner. This approach allows us to more explicitly discern the contribution of each training stage to the overall performance. The experimental results are presented in Figure 4, where $stage_{1}$, $stage_{2}$, and $stage_{3}$ refer to the training phases involving simple, common, and hard samples, respectively. We observed that the performance of both the retriever and the generator steadily improves as the training stages progress. Furthermore, the removal of the curriculum learning strategy resulted in a significant performance decline, indicating that neglecting the difficulty order of training samples leads to sub-optimal training outcomes.
5 Conclusion
In this study, we introduce CL-RAG, a curriculum learning-based training framework designed for the RAG system. To the best of our knowledge, it’s one of the first time that the idea of human imitation learning is integrated with RAG training, which can effectively enhance the generalization and stability of the RAG system. Experiments on four open-domain question answering datasets provide substantial evidence of the framework’s effectiveness. Additionally, separate experiments conducted on the retriever and the generator have demonstrated the significant enhancements our method brings to each individual part.
Limitations
Despite its effectiveness, CL-RAG still has certain limitations. First, the LLM preference measurement based on the probability or ranking of decoding correct answers may have advantages in open-domain question answering data, but it may be unstable in long text generation. Second, our framework only involves a single iteration of training between the RALM and the retriever, that is, training the RALM first and then the retriever. More refined iterative methods, such as iteration after each difficulty level training stage, still need further exploration.
Appendix A Datasets Details
@@
Table 4: Details of the datasets we used.
Table 4 shows details of the datasets we used, where the data volumes for RALM and Retriever represent the amounts used for training each respective component.
Appendix B Inference Prompt
During the inference stage of all models, we used the same prompt to ensure fairness. As shown in Table 5, paragraph represents the retrieved document, and instruction represents the query. Finally, we extracted the content following "Answer:" in the model’s response as the final answer.
@@ Appendix B Inference Prompt
System Prompt: Answer the following questions with two to three words. Your answer must be
formatted as follows: Answer: <your answer>
User Prompt: The following contexts will help you answer the question.
@@
Table 5: Prompt for LLM inference.
Appendix C Complete Results of Robustness Test
Complete results are shown in Table 6.

### v2

@@ III-C Data Augmentation and Curriculum Learning for Retriever
$\mathcal{D}_{rerank}=\{d_{1},d_{2},...,d_{n}\}.$ (3)
Here, $d_{1}$ represents the document most beneficial for the LLM to generate the correct answer, i.e., the document that yields both the maximum rank improvement and the highest probability gain for the correct answer when incorporated with question.
@@
Next, we construct data with different difficulty levels for curriculum learning, aiming to first train the retriever to distinguish documents with obvious quality gaps and then train it to distinguish documents with slightly quality differences. Specifically, we divide $\mathcal{D}_{rerank}$ into three groups $\{\mathcal{D}_{1},\mathcal{D}_{2},\mathcal{D}_{3}\}$ based on the ranking order from RALM. These groups contain documents in the ranges $d_{1},d_{n_{1}}]$, $(d_{n_{1}},d_{n_{2}}]$, and $(d_{n_{2}},d_{n}]$, respectively, representing good documents, sub-optimal documents, and hard negative documents for answering questions.
For the retriever, we expect it to distinguish the difference between good and bad documents, and to assign higher scores to good documents and lower scores to bad documents. When the document quality in the training data varies significantly (with obvious good and bad documents), the retriever can more easily distinguish the quality differences among documents and assign scores with larger gaps, making training easier. When the document quality in the training data becomes more consistent (either all relatively good or all relatively poor), the retriever struggles to distinguish between these documents, assigns very similar scores to them, and training becomes more difficult. Therefore, we define training data with a larger ranking span between different documents in the sampled data as easy data—i.e., the simplest data is sampled across $d_{1},d_{n_{1}}]$, $(d_{n_{1}},d_{n_{2}}]$, and $(d_{n_{2}},d_{n}]$. The smaller the ranking span, the more difficult the data is for the retriever—e.g., data sampled only within $d_{1},d_{n_{1}}]$ is the most difficult.
TABLE II: Divergence in LLM-based Document Ranking Across Retriever Training Stages
[…]
Easy 1 8.39 19
Common 2 4.90 14
Hard 3 2.00 4
To construct our stage-wise training data, we sample $k_{1}$, $k_{2}$, and $k_{3}$ documents from each of the three groups, where $k_{1}+k_{2}+k_{3}=k$. Across stages, we increase $k_{1}$ while decreasing $k_{2}$ and $k_{3}$, thereby narrowing the quality gap between the sampled documents. Table II presents the average ranking divergence of documents across different training phases. Larger values indicate more distinct document rankings, demonstrating that these documents can be more easily distinguished by the retriever. Conversely, smaller ranking differences imply that the documents are more difficult to differentiate in terms of quality. We have also designed a tiered loss function to better fit our approach:
$\mathcal{L}(q,\mathcal{D}_{k})=\\ -\sum_{d_{i,j}\in{\mathcal{D}_{k}}}\sum_{j>i}\frac{j-i}{n-1}log(\frac{e^{s_{i}}}{e^{s_{i}}+e^{s_{j}}}),$ (4)
@@
where $\mathcal{D}_{k}$ denotes a subset of $k$ documents selected from $\mathcal{D}$, $n$ represents the total number of documents involved in the ranking, and $i$,$j$ are the ranks of the documents, $s_{i}$ denotes the similarity score between the query $q$ and the document $d_{i}$. This loss function implies that: (1) the difference in similarity scores should increase with ranking distance, and (2) higher-ranked documents must maintain larger similarity scores. We start by distinguishing documents with obvious quality differences and gradually move to distinguishing documents with slight quality differences. This stage-by-stage approach helps bridge the gap between the original retriever and the retriever aligned with RALM preferences.
IV Experimental Setting
IV-A Datasets
TABLE III: Details of the datasets we used.
@@ IV-A Datasets
2Wiki - - - 12576
IV-A1 Open-domain QA
We considered five open-domain QA datasets: single-hop question datasets Natural Questions (NQ) [47], TriviaQA [48], and PopQA [49], as well as the multi-hop question dataset HotpotQA [50] and 2WikiMultiHopQA (2Wiki)[51]. For each dataset, we employed Contriever-MSMARCO [52] as the retriever. Following previous settings [31], for NQ, TriviaQA, HotpotQA, and 2Wiki, we retrieved the corresponding documents from the 2018 Wikipedia corpus provided by [10]. For PopQA, we retrieved the documents from the 2020 Wikipedia corpus. Noting that PopQA only contains a test set, which we split into two non-overlapping parts for training and testing, respectively. We utilized datasets from the KILT benchmark [53] and open-source platform Hugging Face [^1] The datasets for NQ, TriviaQA, and HotpotQA were sourced from the KILT benchmark, while PopQA and 2Wiki datasets were obtained from the authors’ publicly released data.. We sampled training data from the training sets of NQ, TriviaQA, HotpotQA, and PopQA, while reserving 2Wiki exclusively for testing purposes without including it in the training process. Table III details the composition of our training data, where RALM and Retriever denote the randomly sampled subsets from the training set for generator and retriever training, respectively.
IV-A2 Robustness Test
IV-B Evaluation Metrics
We employ Exact Match (EM) and F1 score to evaluate the effectiveness of final answer generation. Specifically, EM assesses whether the LLM-generated answers are identical to the correct answers. Meanwhile, the F1 score integrates Precision and Recall to measure the accuracy and coverage of the LLM in generating answers.
For assessing the quality of documents retrieved by the retriever, we employ Recall@k (R@k) as the evaluation metric. This metric indicates the proportion of top-k retrieved documents that contain the correct answer. Specifically, $R@5=50\%$ means that for $50\%$ of the questions, their correct answers are included within the top-5 retrieved documents.
IV-C Baselines
We utilized several training methods for RALM and Retriever as baselines, and firstly evaluated the performance of all the RALM training methods: (1) RAAT[14]: Enhance the robustness of RALMs through adversarial training. (2) RALMtop5: The most basic approach in RALM training by directly employing the toptop-55 retrieved documents. (3) RALMgolden[13]: Add the golden document to the retrieved documents to train the RALM. (4) RetRobust[12]: Randomly inject irrelevant and counterfactual documents into the retrieved documents to expose the model to diverse types of noise during training.
Then, we continued to assess the performance of retrievers on the best-performing RALM: (1) Replug[11]: Minimize the KL divergence between the score of retriever’s and the model’s preference distribution to train a retriever that better aligns with the LLM. (2) RA-DIT[13]: Combine the method of finetuning the retriever using KL divergence and finetuning the RALM with retrieved documents that include golden documents as context. (3) LLM-Embedder[21]: Define the quality of document $d$ based on how much it improves the correct answer’s ranking in LLM’s response and proposes a fine-grained hierarchical loss function for training.
TABLE IV: Prompt for LLM inference.
@@ IV-C Baselines
System Prompt: Answer the following questions with two to three
words. Your answer must be formatted as follows:
Answer: <<your answer>>
User Prompt: The following contexts will help you answer the
question.
@@
TABLE V: Experimental results for EM and F1 scores(%) on four open-domain QA datasets compared with multiple baselines in RALM or Retriever training. The underlined and boldface items indicate the best results under each setting. $RALM_{CL}$” refers to the RALM trained with our DACL strategy.